# Towards Data Contamination Detection for Modern Large Language Models: Limitations, Inconsistencies, and Oracle Challenges

### v1

@@ Abstract
As large language models achieve increasingly impressive results, questions arise about whether such performance is from generalizability or mere data memorization. Thus, numerous data contamination detection methods have been proposed. However, these approaches are often validated with traditional benchmarks and early-stage LLMs, leaving uncertainty about their effectiveness when evaluating state-of-the-art LLMs on the contamination of more challenging benchmarks. To address this gap and provide a dual investigation of SOTA LLM contamination status and detection method robustness, we evaluate five contamination detection approaches with four state-of-the-art LLMs across eight challenging datasets often used in modern LLM evaluation. Our analysis reveals that (1) Current methods have non-trivial limitations in their assumptions and practical applications; (2) Notable difficulties exist in detecting contamination introduced during instruction fine-tuning with answer augmentation; and (3) Limited consistencies between SOTA contamination detection techniques. These findings highlight the complexity of contamination detection in advanced LLMs and the urgent need for further research on robust and generalizable contamination evaluation. Our code is available at https://github.com/vsamuel2003/data-contamination
@@
Vinay Samuel1∗  Yue Zhou2∗  Henry Peng Zou2 1Carnegie Mellon University 2University of Illinois Chicago vsamuel@andrew.cmu.edu, {yzhou232,pzou3}@uic.edu
**footnotetext: Authors contributed equally to this work.
@@ 1 Introduction
While large language models (LLMs) consistently achieve higher state-of-the-art results across various benchmarks (Rae et al., 2022; Srivastava et al., 2023; Liang et al., 2023; Zou et al., 2024), the lack of curation in and the limited disclosure of massive training datasets raise a critical question: Does the model performance arise from model generalizability or mere memorization? Furthermore were the test sets possibly contaminated without notice? These questions have become crucial in accurately gauging LLMs’ performance and has led to a critical area of research: detecting data contamination in LLMs.
Recent work in detecting data contamination in LLMs has primarily focused on detecting contamination through validating the log probability of the data in the datasets (Oren et al., 2024; Shi et al., 2023) or determining contamination through prompting-based approaches (Golchin and Surdeanu, 2024b, a). Additional studies have been aimed at understanding the different types of contamination such as differentiating between training/development split contamination and testing set contamination (Sainz et al., 2023) and contamination occurring in the pertaining phase versus the supervised fine-tuning stage (Jacovi et al., 2023).
However, existing research in this area has several limitations. Firstly, most methods show their effectiveness on traditional benchmarks, which are likely overexposed online and to LLMs. In contrast, challenging datasets that test the limits of LLM capabilities are neglected. These newer benchmarks are often more complex with novel formats. Secondly, the tested LLMs are often early-staged ones, such as GPT-J, while the rapid pace of LLM development has left a gap in understanding contamination in the latest models. Thirdly, previous research predominantly focuses on (possibly unintentional) contamination occurring during pretraining, where models are exposed to data in its original form. However, it overlooks contamination during instruction fine-tuning, where original data is subtly modified by, for instance, answer augmentation with chain-of-thought reasoning steps. Such variations may cause difficulties in contamination detection yet are rarely considered when evaluating the effectiveness of contamination detection methods. Lastly, while various detection methods have demonstrated effectiveness (with a particular set of datasets and LLMs of their choice), there is a lack of a comprehensive cross-comparison to assess the consistency and reliability of these techniques, particularly for more recent LLMs, benchmarks, and training paradigms.
@@ 3 Benchmark Datasets
There exists a significant disparity between the benchmarks commonly used in data contamination research and those employed to evaluate state-of-the-art (SOTA) LLM capabilities. Table 4 illustrates this mismatch, showing that while many contamination detection methods are validated using traditional benchmarks, modern LLM evaluations focus on more challenging tasks, such as mathematical reasoning and code generation. This observation raises intriguing questions: Could the performance gains of SOTA LLMs on these newer, more challenging benchmarks be attributed, in part, to data contamination? What insights do SOTA data contamination detection methods provide when applied to these essential modern benchmarks?
@@ 5.1 Implementation Details
We used $k=20$ as recommended in the paper. The authors used their proposed WikiMIA dataset as ground truth and reported AUC in the original paper. However, the threshold for determining contamination is not provided, and in real-world settings, the ground truth is unavailable. Thus, we used the mean and standard deviation for each split directly for the contamination indicator, as the authors claim Min-KPorb is most informative compared with other probability-based metrics.
@@
We used the exact implementation without changes, including only ten instances. GPT-4 was employed to determine near/exact matches, with one exact match and two near matches used as the threshold for contamination detection. The temperature was set to 0 for inference.
@@
To investigate the challenges of detecting data contamination during the fine-tuning stage, we developed an oracle setup that mimics real-world instruction fine-tuning scenarios. We hypothesize that contamination occurring during fine-tuning is significantly more difficult to detect compared to pretraining contamination for several key reasons: (1) Exposure Frequency: During fine-tuning, the model typically encounters data for only a few epochs (usually 1-3), whereas in pretraining, data chunks are often seen repeatedly due to sliding window approaches ([starting position: starting position + window size]). (2) Data Modification: Modern fine-tuning techniques, particularly instruction fine-tuning, often incorporate more complex answers, such as chain-of-thought reasoning, to enhance the LLM’s analytical capabilities. This process modifies the original data, preserving the questions while replacing answers with elaborate solutions. To simulate these conditions, we fine-tuned LLaMA-2-70B-chat using varying proportions of the six newer benchmarks in our study. Our goal was to observe how different contamination detection metrics respond to varying levels of data exposure during fine-tuning. For each data point, we replaced the original answer with a chain-of-thought solution generated by LLaMA-2-70B, mimicking the data augmentation often used in instruction fine-tuning. To maintain some semblance of the original data structure, we packaged four examples as one training instance, preserving the local order as it appears in the original dataset. This allows to test whether detection methods can identify contamination when local order information is partially retained. The fine-tuning process used a learning rate of 8e-6 and ran for three epochs, aligning with typical fine-tuning practices. By varying the proportion of benchmark data used in this process, we aimed to create a controlled environment where the degree of contamination is known, allowing us to evaluate the sensitivity and reliability of various detection methods in identifying fine-tuning stage contamination.
@@ 5.2 Main Results
Table 1 shows the results of the five contamination detection methods on eight benchmarks with GPT-4, Claude-3, and LLaMA-3. Results for the Min-Kand Canonical Order methods are unavailable for GPT-4 and Claude-3, as these approaches require access to model parameters. We can observe that: (1) Perturbation accuracy is unusually high across most benchmarks for all three models. This consistent pattern suggests potential inflation of results, possibly due to the models recognizing the perturbations themselves rather than indicating contamination or memorization. (2) Both our proposed method and the token overlap approach provide significant evidence that the IMDB dataset may be contaminated for all three models. (3) The token overlap method yields conflicting results depending on the specific metric. Based on p-values, it suggests contamination in HumanEval for Claude-3 and in BBH for GPT-4. However, no contamination is detected for any of the newer benchmarks when considering exact matches and near matches. (4) For LLaMA-3, both Min-Kand Canonical Order P-value show possible contamination on BBH. However, in general, there is no clear agreement among the methods regarding contamination in the newer benchmarks. Notably, the GPT-4 report indicates that DROP and HumanEval are approximately 21-25% contaminated. However, our results show that none of the methods detected contamination in DROP, while only our method partially detected contamination in HumanEval (accuracy = 0.36).
@@
Table 2 presents the oracle results, comparing the metric values of the five detection methods before (upper half) and after (lower half) instruction fine-tuning. The Portion column indicates the percentage of data used for fine-tuning, ranging from 0% (no contamination) to 100% (entire dataset split used). Ideally, a robust detection method should (1) consistently reflect the portion of contamination through its metric values and (2) not be influenced by the training set contamination when the test set remains uncontaminated. For instance, probability-based approaches might be susceptible to false positives in test set contamination detection when the training set is contaminated due to the similarity in distribution between training and testing data. Therefore, post-fine-tuning results should show increased accuracy/mean probability values and decreased p-values for non-zero contamination portions in an ideal scenario. However, our observations reveal significant challenges in detecting fine-tuning contamination: (1) None of the metrics demonstrate a consistent trend that aligns with the varying portions of data contamination. This lack of correlation suggests that current methods may not be sensitive enough to detect or quantify the degree of contamination introduced during such a fine-tuning paradigm. (2) Surprisingly, all Min-Kprobability values decrease after fine-tuning. This counterintuitive result is particularly noteworthy given that the original questions are preserved in the fine-tuning process, even though we used instruction fine-tuning with chain-of-thought solutions. These findings highlight the complexity of detecting contamination in instruction-tuned models and suggest that existing methods, originally designed for pretraining contamination detection, may not be directly applicable or reliable in fine-tuning scenarios.
@@ 6 Conclusions and Future Work
Our study offers valuable insights into data contamination detection in state-of-the-art LLMs over diverse benchmarks and models. However, the detection methods we examined, including ours, showed limited effectiveness in accurately reflecting the known contamination in our oracle setup. This unexpected result points to the complexity of contamination mechanisms in instruction-tuned models and opens new avenues for research into how fine-tuning affects the detectability of contamination.
@@ A.2 Local Order Quiz
Table 4: The dataset mismatch between the challenging benchmarks frequently evaluated by SOTA LLMs and the ones used for validation of contamination detection methods.

### v2

@@ Abstract
As large language models achieve increasingly impressive results, questions arise about whether such performance is from generalizability or mere data memorization. Thus, numerous data contamination detection methods have been proposed. However, these approaches are often validated with traditional benchmarks and early-stage language models, leaving uncertainty about their effectiveness when evaluating state-of-the-art ones with more challenging benchmarks. To address this gap and provide a dual investigation of the contamination status of state-of-the-art large language models and detection method robustness, we evaluate five contamination detection approaches with four state-of-the-art large language models across eight challenging and prevailing datasets. Our analysis reveals that (1) Current methods have non-trivial limitations in their assumptions and practical applications; (2) Notable difficulties exist in detecting contamination introduced during instruction fine-tuning with answer augmentation; and (3) Limited consistencies between state-of-the-art contamination detection techniques. These findings highlight the complexity of contamination detection in advanced language models and the urgent need for further research on robust and generalizable contamination evaluation. Our code is available at https://github.com/vsamuel2003/data-contamination
@@
Vinay Samuel1thanks: Authors contributed equally to this work.thanks: Correspondence: vsamuel@andrew.cmu.edu, Yue Zhou211footnotemark: 1, Henry Peng Zou2 1Carnegie Mellon University, 2University of Illinois Chicago
@@ 1 Introduction
While large language models (LLMs) consistently achieve higher state-of-the-art results across various benchmarks (Rae et al., 2022; Srivastava et al., 2023; Liang et al., 2023; Zou et al., 2024), the lack of curation in and the limited disclosure of massive training datasets raise a critical question: Does the model performance arise from model generalizability or mere memorization? Furthermore, were the test sets possibly contaminated without notice? These questions have become crucial in accurately gauging LLMs’ performance and has led to a critical area of research: detecting data contamination in LLMs.
Data contamination occurs when test or evaluation data is exposed to the model during its training phases (either pre-training or fine-tuning). This can lead to artificially inflated performance metrics through memorization rather than true generalization. Recent work in detecting data contamination in LLMs has primarily focused on detecting contamination through validating the log probability of the data in the datasets (Oren et al., 2024; Shi et al., 2023) or determining contamination through prompting-based approaches (Golchin and Surdeanu, 2024b, a). Additional studies have been aimed at understanding the different types of contamination, such as differentiating between training or development split contamination and testing set contamination (Sainz et al., 2023) and contamination occurring in the pertaining phase versus the supervised fine-tuning stage (Jacovi et al., 2023).
However, existing research in this area has several limitations. Firstly, most methods are shown to be effective on traditional benchmarks, which are likely overexposed online and to LLMs. In contrast, challenging datasets that test the limits of LLM capabilities are neglected. These newer benchmarks are often more complex with novel formats. Secondly, the tested LLMs are often early-staged ones, such as GPT-J, while the rapid pace of LLM development has left a gap in understanding contamination in the latest models. Thirdly, previous research predominantly focuses on (possibly unintentional) contamination occurring during pretraining, where models are exposed to data in its original form. However, it overlooks contamination during instruction fine-tuning, where original data is subtly modified by, for instance, answer augmentation with chain-of-thought reasoning steps. Such variations may cause difficulties in contamination detection yet are rarely considered when evaluating the effectiveness of contamination detection methods. Lastly, while various detection methods have demonstrated effectiveness (with a particular set of datasets and LLMs of their choice), there is a lack of a comprehensive cross-comparison to assess the consistency and reliability of these techniques, particularly for more recent LLMs, benchmarks, and training paradigms.
@@ 2 Related Work
Besides, recent work by Yao et al. (2024) introduces a generalization-based approach, highlighting how cross-lingual contamination can bypass traditional detection methods by inflating performance when benchmarks are translated into other languages. This study shows that models contaminated in this way fail to generalize effectively when answer options are manipulated, offering a novel angle for detecting deeply concealed contamination.
@@ 3 Benchmark Datasets
There exists a significant disparity between the benchmarks commonly used in data contamination research and those employed to evaluate state-of-the-art (SOTA) LLM capabilities. Table A1 in the Appendix illustrates this mismatch, showing that while many contamination detection methods are validated using traditional benchmarks, modern LLM evaluations focus on more challenging tasks, such as mathematical reasoning and code generation. This observation raises intriguing questions: Could the performance gains of SOTA LLMs on these newer, more challenging benchmarks be attributed, in part, to data contamination? What insights do SOTA data contamination detection methods provide when applied to these essential modern benchmarks?
@@ 5.1 Implementation Details
We used $k=20$ as recommended in the paper. The authors used their proposed WikiMIA dataset as ground truth and reported AUC in the original paper. However, the threshold for determining contamination is not provided, and in real-world settings, the ground truth is unavailable. Thus, we used the mean and standard deviation for each split directly for the contamination indicator, as the authors claim Min-KProb is most informative compared with other probability-based metrics.
@@
We used the exact implementation without changes, including only ten instances. GPT-4 was employed to determine near/exact matches, with one exact match and two near matches used as the threshold for contamination detection. The temperature was set to 0 for inference. The exact ICL prompt used for GPT-4 evaluation is shown in Figure A.2 in the Appendix.
@@
To investigate the challenges of detecting data contamination during the fine-tuning stage, we developed an oracle setup that mimics real-world instruction fine-tuning scenarios. We hypothesize that contamination occurring during fine-tuning is significantly more difficult to detect compared to pretraining contamination for several key reasons: (1) Exposure Frequency: During fine-tuning, the model typically encounters data for only a few epochs (usually 1-3), whereas in pretraining, data chunks are often seen repeatedly due to sliding window approaches ([starting position: starting position + window size]). (2) Data Modification: Modern fine-tuning techniques, particularly instruction fine-tuning, often incorporate more complex answers, such as chain-of-thought reasoning, to enhance the LLM’s analytical capabilities. This process modifies the original data, preserving the questions while replacing answers with elaborate solutions. To simulate these conditions, we fine-tuned LLaMA-2-70B-chat using full fine-tuning with varying proportions of the six newer benchmarks in our study. Our goal was to observe how different contamination detection metrics respond to varying levels of data exposure during fine-tuning. For each data point, we replaced the original answer with a chain-of-thought solution generated by LLaMA-2-70B, mimicking the data augmentation often used in instruction fine-tuning. To maintain some semblance of the original data structure, we packaged four examples as one training instance, preserving the local order as it appears in the original dataset. This allows to test whether detection methods can identify contamination when local order information is partially retained. The fine-tuning process used a learning rate of 8e-6 and ran for three epochs, aligning with typical fine-tuning practices. By varying the proportion of benchmark data used in this process, we aimed to create a controlled environment where the degree of contamination is known, allowing us to evaluate the sensitivity and reliability of various detection methods in identifying fine-tuning stage contamination.
@@ 5.2 Main Results
Table 1 shows the results of the five contamination detection methods on eight benchmarks with GPT-4 (gpt-4-0613), Claude-3 Haiku, and LLaMA-3-70b. Results for the Min-Kand Canonical Order methods are unavailable for GPT-4 and Claude-3, as these approaches require access to model parameters. We can observe that: (1) Perturbation accuracy is unusually high across most benchmarks for all three models. This consistent pattern suggests potential inflation of results, possibly due to the models recognizing the perturbations themselves rather than indicating contamination or memorization. (2) Both our proposed method and the token overlap approach provide significant evidence that the IMDB dataset may be contaminated for all three models. (3) The token overlap method yields conflicting results depending on the specific metric. Based on p-values, it suggests contamination in HumanEval for Claude-3 and in BBH for GPT-4. However, no contamination is detected for any of the newer benchmarks when considering exact matches and near matches. (4) For LLaMA-3, both Min-Kand Canonical Order P-value show possible contamination on BBH. However, in general, there is no clear agreement among the methods regarding contamination in the newer benchmarks. Notably, the GPT-4 report indicates that DROP and HumanEval are approximately 21-25% contaminated. However, our results show that none of the methods detected contamination in DROP, while only our method partially detected contamination in HumanEval (accuracy = 0.36).
@@
Table 2 presents the oracle results, comparing the metric values of the five detection methods before (upper half) and after (lower half) instruction fine-tuning. The Portion column indicates the percentage of data used for fine-tuning, ranging from 0% (no contamination) to 100% (entire dataset split used). Ideally, a robust detection method should (1) consistently reflect the portion of contamination through its metric values and (2) not be influenced by the training set contamination when the test set remains uncontaminated. For instance, probability-based approaches might be susceptible to false positives in test set contamination detection when the training set is contaminated due to the similarity in distribution between training and testing data. Therefore, post-fine-tuning results should show increased accuracy or mean probability values and decreased p-values for non-zero contamination portions in an ideal scenario. However, our observations reveal significant challenges in detecting fine-tuning contamination: (1) None of the metrics demonstrate a consistent trend that aligns with the varying portions of data contamination. This lack of correlation suggests that current methods may not be sensitive enough to detect or quantify the degree of contamination introduced during such a fine-tuning paradigm. (2) Surprisingly, all Min-Kprobability values decrease after fine-tuning. This counterintuitive result is particularly noteworthy given that the original questions are preserved in the fine-tuning process, even though we used instruction fine-tuning with chain-of-thought solutions. These findings highlight the complexity of detecting contamination in instruction-tuned models and suggest that existing methods, originally designed for pretraining contamination detection, may not be directly applicable or reliable in fine-tuning scenarios.
@@ 6 Conclusions and Future Work
While we discovered significant disagreement between different contamination detection methods, especially compared to oracle tests, our investigation did not uncover the underlying reasons for these discrepancies. Additionally, our proposed prompt-based contamination detection method showed limited effectiveness, highlighting the inherent challenges in using prompting for this task.
@@ A.2 Local Order Quiz
GPT-4 ICL prompt used for Token Overlap INSTRUCTION:
You are provided with a reference text and a candidate text. The candidate text is a generated replica of the reference text. Your task is to determine if the candidate text is an exact or near-exact match of the reference text.
Near-exact match candidates must PRECISELY match the reference candidates in terms of sentence structure, overlap, and contextual similarity. Respond only with ’Yes’ or ’No’. Example 1:
REFERENCE TEXT:
The cat waited at the top. CANDIDATE TEXT: The cat waited at the top. ANSWER: Yes (exact match) Example 2:
REFERENCE TEXT:
icy surface of Jupiter’s largest moon, Ganymede. These irregular masses may be rock formations, supported by Ganymede’s icy shell for billions of years. CANDIDATE TEXT: icy surface of Jupiter’s largest moon, Ganymede. These irregular masses may be rock formations, supported by Ganymede’s icy shell for billions of years. This discovery supports the theory that Ganymede has a subsurface ocean. Scientists used gravity data from NASA’s Galileo spacecraft to create a geophysical model of the interior of Ganymede. ANSWER: Yes (near-exact match) Example 3:
REFERENCE TEXT:
50th Anniversary of Normandy Landings lasts a year. CANDIDATE TEXT: The 50th anniversary celebration of the first Normandy landing will last a year. ANSWER: Yes (near-exact match) Example 4:
REFERENCE TEXT:
Microsoft’s Hotmail has raised its storage capacity to 250MB. CANDIDATE TEXT: Microsoft has increased the storage capacity of its Hotmail e-mail service to 250MB. ANSWER: Yes (near-exact match) Example 5:
REFERENCE TEXT:
{reference_text} CANDIDATE TEXT: {candidate_text} ANSWER:
@@
Table A1: The dataset mismatch between the challenging benchmarks frequently evaluated by SOTA LLMs and the ones used for validation of contamination detection methods.